# ADNF-Clustering: An Adaptive and Dynamic Neuro-Fuzzy Clustering for Leukemia Prediction

### v1

@@ Abstract
Index Terms:
Adaptive clustering, dynamic fuzzy clustering, leukemia, medical image analysis, neuro‐fuzzy systems, streaming algorithms.
@@ I-A Challenges and Unmet Needs
Despite the fact that fuzzy clustering is very effective in leukemia images, it does not fit well in dynamic and adaptive situations, in which the images are continuously evolving. We created a new algorithm, called Adaptive and Dynamic Neuro Fuzzy Clustering that goes under the acronymous of ADNF clustering. In particular, the proposed methodology holds immediate potential for integration within the INFANT project, a regional initiative aimed at enhancing pediatric oncology care in Campania. The project’s hub-and-spoke network, centered around the Santobono-Pausilipon hospital, provides an ideal framework for future refinement and validation of the model using consolidated local patient data, ultimately contributing to improved diagnostic capabilities across various healthcare contexts. The innovation of this method consists in the implementation of the dynamic part within the fuzzy clustering. Mainly, the fuzzy is useful to capture the transitions between cells, but it does that in a static way. In our project, the fuzzy system acts like an adaptive clustering mechanism that is capable of changing its hyperparameters and thresholds over time. In particular, we extracted the features of the cells using a Convolutional Neural Network (CNN). From this, we obtain the first input for every image: a feature vector $X\in\mathbb{R}^{n}$. At this point, the fuzzy c‑means algorithm can be applied to the batch to obtain a set of micro‑clusters. For each cluster, the centroid $u_{i}$ is computed and the membership degrees $u_{ij}$ are initially assigned using a fixed value of m. All initial statistics are then calculated, such as number of points, density, timestamp, dynamic threshold $\varepsilon_{i}$. Then we extract a new vector. At this stage, our dynamic clustering comes into play. Whenever a new data point arrives, we must calculate the distance from the new point $X$ to the centroid of each micro‑cluster $u_{i}$. Then we assign the point with a certain fuzzy deegree, updating the centroid with the decay parameter $\lambda_{t}^{j}$. After this, it’s always important to update the density, timestamp and number of points.Before getting into the merging and splitting of the final clusters, we calculated the local fuzzy entropy and the fuzzy transactional index, alias FTI.
@@ II-A Limitations of Supervised Approaches
Despite impressive results, most clinical CNN systems rely on supervised learning, which demands large volumes of expert‐annotated images. High‐quality labeling in medicine is expensive, time‑consuming, and subject to inter‐observer variability: experts may disagree, and annotations can be probabilistic or context‑dependent [19]. Label noise introduced by these inconsistencies can degrade model accuracy and generalization. Moreover, models trained on data from a single institution often suffer from dataset shift when deployed elsewhere, due to differences in imaging protocols, hardware, and patient populations [20]. While techniques such as data augmentation, transfer learning, and domain adaptation mitigate these issues [21], they do not eliminate the fundamental dependency on labeled data.
II-B Unsupervised Feature Extraction
To reduce reliance on labels, unsupervised and self‑supervised methods have gained traction. A popular strategy is to repurpose pretrained CNNs as fixed feature extractors: by forwarding raw medical images through an ImageNet‑trained or domain‑specific network (with its classification head removed), one obtains high‑dimensional embeddings that capture semantic content without requiring new annotations [22]. These embeddings discard irrelevant pixel‑level noise while preserving diagnostically salient information, enabling subsequent analysis in a lower‑dimensional feature space.
II-C Clustering in Medical Image Analysis
Clustering these CNN‑extracted features uncovers latent structure in large, unlabeled datasets [23]. In medical contexts, clusters may correspond to disease subtypes, anatomical variants, progression stages, or technical artifacts. Clustering aids dataset curation (detecting outliers, redundant samples, or potential mislabels) and can guide annotation by selecting representative examples for expert review [7].
II-D Challenges of Traditional Clustering
Standard clustering algorithms (e.g., K‑Means [24], Gaussian Mixture Models [25], hierarchical clustering [26]) face several obstacles in medical imaging: fixed partitioning assumes hard boundaries, static structure requires retraining to incorporate new data, and lack of uncertainty modeling obscures mixed or transitional cases.
II-E Fuzzy and Dynamic Clustering Paradigms
@@
Fuzzy clustering allows partial membership across clusters [3], aligning with the soft boundaries of early‑stage or heterogeneous pathologies. Yet classical fuzzy methods remain static. Dynamic clustering algorithms update centroids and memberships incrementally as new samples arrive [9], and combining this with fuzzy logic provides both adaptability and uncertainty quantification. Incorporating entropy‑based criteria [8] permits the algorithm to monitor its own uncertainty: high internal entropy can trigger a split, while low inter‑cluster distances may prompt merging, balancing stability against plasticity.
II-F Towards Adaptive Neuro‑Fuzzy Clustering
@@
In this work, we propose a Dynamic Neuro‑Fuzzy Clustering framework tailored to medical image analysis. We extract semantic embeddings via a pretrained CNN and feed them into an adaptive clustering pipeline that:
@@
    Produces initial soft partitions via Fuzzy C‑Means;
@@
    Updates micro‑clusters online with a dynamic learning rate that reflects recent density and positional changes;
@@
    Tunes cluster fuzziness through a Fuzzy Temporal Index (FTI) measuring entropy evolution;
@@
    Refines topology by density‑weighted merging and entropy‑guided splitting.
This approach requires no labels, adapts continuously to incoming images, and models uncertainty explicitly—qualities crucial for scalable, flexible, and clinically meaningful analysis of large‑scale medical imaging datasets.
III Adaptive Dynamic Neuro‑Fuzzy Clustering
Adaptive Dynamic Neuro‑Fuzzy Clustering is a streaming‑capable hybrid algorithm that seamlessly marries the soft‑assignment strengths of Fuzzy C-Means with a suite of dynamic, data-driven mechanisms-micro-cluster updates, density-weighted merging, and entropy-guided splitting. Its core ambition is to yield high-fidelity models in contexts where observations arrive continuously or where the underlying distribution drifts over time, as is typical in biomedical imaging or real‑time monitoring scenarios.
@@ III Adaptive Dynamic Neuro‑Fuzzy Clustering
Let $\mathcal{X}=\{x_{j}\}_{j=1}^{N}\subset\mathbb{R}^{d}$ denote our set of $d$‑dimensional feature vectors, extracted via a pretrained convolutional backbone. We choose an initial cluster count $c$ and a base fuzziness $m_{0}>1$. Each micro‑cluster $MC_{i}$ encapsulates:
@@ IV Implementation
We used microscopy images from the C‑NMC leukemia dataset [14], which contains 15,135 RGB segmented cell images from 118 patients. The dataset is divided into two classes: normal cells and leukemic blasts. These cells have been extracted from real-world blood smear images and exhibit characteristics such as staining noise and illumination variation, although most artifacts were corrected during acquisition. All images have been resized to $64\times 64$ pixels regardless of their original dimensions. This ensures a consistent input size for the model while also reducing computational complexity. Subsequently, each pixel value was normalized by dividing by 255, bringing the data into the [0,1] range.
@@
5:pca_features $\leftarrow$ PCA(n_c=10).fit_transform(features_flat)
@@ IV-A ADNF Clustering Pipeline
2:$(\texttt{centroids},U,\dots)\leftarrow\texttt{cmeans}(\mathbf{X}^{\top},c,m_{0})$
@@
  • Split high-entropy clusters via DBSCAN on points with local entropy $>\tau_{s}$.
@@ V Conclusions
Adaptive and Dynamic Neuro–Fuzzy Clustering overcomes key limitations of static clustering approaches in leukemia image analysis by combining CNN‐based feature extraction with an online fuzzy clustering engine designed for streaming data. Through continuous micro‐cluster updates regulated by a Fuzzy Temporal Index, our approach dynamically adjusts cluster boundaries and fuzziness to reflect evolving cellular patterns while providing quantifiable uncertainty measures. The topology‑refinement steps (density‑weighted merging and entropy‑guided splitting) ensure robustness against both over‑ and under‑segmentation. Empirical evaluation on the C‑NMC leukemia microscopy dataset demonstrates a silhouette score of 0.51, outperforming conventional static baselines in cluster cohesion and separation.it’s important to remember that our clustering is not only fuzzy, but also dynamic. This, combined with the lack of medical consultation or suitable medical datasets, makes this result a major achievement.Furthermore, nowadays the most important Silhouette score of Dynamic Neuro-Fuzzy Clustering non-supervised are approximately between 0.4 and 0.5. The unsupervised, adaptive characteristics of ADNF‑Clustering make it highly suitable for real‑time clinical deployment, offering scalable support for personalized leukemia monitoring within pediatric oncology networks, and aligning perfectly with the INFANT project’s goal of developing deployable, resource‑conscious approaches.

### v2

@@ I-A Challenges and Unmet Needs
Although fuzzy clustering is highly effective for leukemia image analysis, it is not well-suited for dynamic and adaptive scenarios, where images are continuously evolving. To overcome this limitation, we developed a new algorithm for monitoring leukemia at the level of individual patients, allowing us to track the cluster membership of each case. We call this method Adaptive and Dynamic Neuro-Fuzzy Clustering, that goes under the acronymous of ADNF clustering. In particular, the proposed methodology holds immediate potential for integration within the INFANT project, a regional initiative aimed at enhancing pediatric oncology care in Campania. The project’s hub-and-spoke network, centered around the Santobono-Pausilipon hospital, provides an ideal framework for future refinement and validation of the model using consolidated local patient data, ultimately contributing to improved diagnostic capabilities across various healthcare contexts. The innovation of this method consists in the implementation of the dynamic part within the fuzzy clustering. Mainly, the fuzzy is useful to capture the transitions between cells, but it does that in a static way. In our project, the fuzzy system acts like an adaptive clustering mechanism that is capable of changing its hyperparameters and thresholds over time. In particular, we extracted the features of the cells using a Convolutional Neural Network (CNN). From this, we obtain the first input for every image: a feature vector $X\in\mathbb{R}^{n}$. At this point, the fuzzy c‑means algorithm can be applied to the batch to obtain a set of micro‑clusters. For each cluster, the centroid $u_{i}$ is computed and the membership degrees $u_{ij}$ are initially assigned using a fixed value of m. All initial statistics are then calculated, such as number of points, density, timestamp, dynamic threshold $\varepsilon_{i}$. Then we extract a new vector. At this stage, our dynamic clustering comes into play. Whenever a new data point arrives, we must calculate the distance from the new point $X$ to the centroid of each micro‑cluster $u_{i}$. Then we assign the point with a certain fuzzy deegree, updating the centroid with the decay parameter $\lambda_{t}^{j}$. After this, it’s always important to update the density, timestamp and number of points.Before getting into the merging and splitting of the final clusters, we calculated the local fuzzy entropy and the fuzzy transactional index, alias FTI.
@@ II-A Limitations of Supervised Approaches
Despite impressive results, most clinical CNN systems rely on supervised learning, which demands large volumes of expert‐annotated images. High‐quality labeling in medicine is expensive, time‑consuming, and subject to inter‐observer variability: experts may disagree, and annotations can be probabilistic or context‑dependent [19]. Label noise introduced by these inconsistencies can degrade model accuracy and generalization. Moreover, models trained on data from a single institution often suffer from dataset shift when deployed elsewhere, due to differences in imaging protocols, hardware, and patient populations [20]. While techniques such as data augmentation, transfer learning, and domain adaptation mitigate these issues [21], they do not eliminate the fundamental dependency on labeled data. To reduce reliance on labels, unsupervised and self-supervised methods have gained traction. A popular strategy is to repurpose pretrained CNNs as fixed feature extractors: by forwarding raw medical images through an ImageNet-trained or domain-specific network (with its classification head removed), one obtains high-dimensional embeddings that capture semantic content without requiring new annotations [22]. These embeddings discard irrelevant pixel-level noise while preserving diagnostically salient information, enabling subsequent analysis in a lower-dimensional feature space. In this work, we employ EfficientNetV2-M as the backbone feature extractor, as it provides an advantageous trade-off between accuracy and computational efficiency compared to earlier CNN families. Its compound scaling and optimized architecture allow extraction of rich, semantically meaningful representations while keeping inference cost manageable, an aspect particularly relevant in medical imaging pipelines, where both precision and efficiency are critical. After evaluating several candidate architectures, EfficientNetV2-M emerged as the optimal choice, consistently offering the best balance between discriminative power and computational overhead.
II-B Clustering in Medical Image Analysis
@@
Clustering these CNN‑extracted features uncovers latent structure in large, unlabeled datasets [23]. In medical contexts, clusters may correspond to disease subtypes, anatomical variants, progression stages, or technical artifacts. Clustering aids dataset curation (detecting outliers, redundant samples, or potential mislabels) and can guide annotation by selecting representative examples for expert review [7]. Standard clustering algorithms (e.g., K‑Means [24], Gaussian Mixture Models [25], hierarchical clustering [26]) face several obstacles in medical imaging: fixed partitioning assumes hard boundaries, static structure requires retraining to incorporate new data, and lack of uncertainty modeling obscures mixed or transitional cases. Fuzzy clustering allows partial membership across clusters [3], aligning with the soft boundaries of early‑stage or heterogeneous pathologies. Yet classical fuzzy methods remain static. Dynamic clustering algorithms update centroids and memberships incrementally as new samples arrive [9], and combining this with fuzzy logic provides both adaptability and uncertainty quantification. Incorporating entropy‑based criteria [8] permits the algorithm to monitor its own uncertainty: high internal entropy can trigger a split, while low inter‑cluster distances may prompt merging, balancing stability against plasticity.
III Adaptive Dynamic Neuro‑Fuzzy Clustering
@@
In this work, we propose a Dynamic Neuro‑Fuzzy Clustering framework specifically designed for medical image analysis. We first extract semantic embeddings using a pretrained CNN and feed them into an adaptive clustering pipeline that:
@@
    Generates initial soft partitions with Fuzzy C‑Means;
@@
    Updates micro‑clusters online using a dynamic learning rate that accounts for recent density and positional changes;
@@
    Adjusts cluster fuzziness via a Fuzzy Temporal Index (FTI) tracking entropy evolution;
@@
    Refines the cluster topology through density-weighted merging and entropy-guided splitting.
This fully unsupervised approach continuously adapts to incoming images and explicitly models uncertainty, which are key qualities for scalable, flexible, and clinically meaningful analysis of large-scale medical imaging datasets. Adaptive Dynamic Neuro‑Fuzzy Clustering is a streaming-capable hybrid algorithm that combines the soft-assignment benefits of Fuzzy C‑Means with dynamic, data-driven mechanisms such as micro-cluster updates, density-weighted merging, and entropy-guided splitting. Its main goal is to produce high-fidelity models in contexts with continuously arriving data or shifting underlying distributions, as commonly encountered in biomedical imaging and real-time monitoring.
@@
Let $\mathcal{X}=\{x_{j}\}_{j=1}^{N}\subset\mathbb{R}^{d}$ denote our set of dd‑dimensional feature vectors, extracted via a pretrained convolutional backbone. We choose an initial cluster count $c$ and a base fuzziness $m_{0}>1$. Each micro‑cluster $MC_{i}$ encapsulates:
@@ IV Implementation
We used microscopy images from the C‑NMC leukemia dataset [14], which contains 15,135 RGB segmented cell images from 118 patients. The dataset is divided into two classes: normal cells and leukemic blasts. These cells have been extracted from real-world blood smear images and exhibit characteristics such as staining noise and illumination variation, although most artifacts were corrected during acquisition. All images have been resized to $64\times 64$ pixels regardless of their original dimensions. This ensures a consistent input size for the model while also reducing computational complexity. Subsequently, each pixel value was normalized by dividing by 255, bringing the data into the [0,1] range. To further reduce dimensionality and remove redundancy in the extracted features, Principal Component Analysis (PCA) was applied. PCA projects the high-dimensional feature vectors into a lower-dimensional space while preserving the most relevant variance, which improves both computational efficiency and clustering performance. After testing several dimensions, we found that retaining 50 principal components ($\text{PCA}=50$) yielded the best results.
@@
5:pca_features $\leftarrow$ PCA(n_c=50).fit_transform(features_flat)
@@ IV-A ADNF Clustering Pipeline
[…]
@@
  • Split high-entropy clusters via DBSCAN on points with local entropy $\tau_{s}$.
@@ V Conclusions
Adaptive and Dynamic Neuro–Fuzzy Clustering overcomes key limitations of static clustering approaches in leukemia image analysis by combining CNN‐based feature extraction with an online fuzzy clustering engine designed for streaming data. Through continuous micro‐cluster updates regulated by a Fuzzy Temporal Index, our approach dynamically adjusts cluster boundaries and fuzziness to reflect evolving cellular patterns while providing quantifiable uncertainty measures. The topology‑refinement steps (density‑weighted merging and entropy‑guided splitting) ensure robustness against both over‑ and under‑segmentation.
Empirical evaluation on the C-NMC leukemia microscopy dataset demonstrates a Silhouette score of 0.51, outperforming conventional static baselines in terms of both cluster cohesion and separation. This is particularly relevant considering that our approach is not only fuzzy but also dynamic, thereby adapting to the intrinsic variability of biomedical imaging data. The lack of medical consultation or access to fully curated medical datasets makes this result a significant achievement. It is also noteworthy that, in the current state of the art, reported Silhouette scores for Dynamic Neuro-Fuzzy Clustering in unsupervised contexts typically range between 0.4 and 0.5, positioning our method in the upper bound of this empirical spectrum. In addition to the Silhouette analysis, we computed the Davies–Bouldin Index (DBI), which provides an orthogonal perspective by quantifying intra-cluster similarity relative to inter-cluster separation. Our Dynamic Neuro-Fuzzy approach achieved a DBI of 0.61, a value that indicates strong compactness and clear separability among clusters. Since lower DBI values reflect better clustering quality (with values below 1 generally considered robust in biomedical imaging contexts), this further corroborates the effectiveness of the proposed methodology. Taken together, the combination of a high Silhouette score (0.51) and a low Davies–Bouldin Index (0.61) highlights the stability and reliability of our clustering framework. These complementary metrics reinforce that the introduced dynamic neuro-fuzzy paradigm is not only competitive but also particularly well-suited for complex and high-variability datasets such as leukemia cell microscopy images.